# Invariant subspace perturbations related to defective eigenvalues of $\Delta$-Hermitian and Hamiltonian matrices

### v1

@@ 4.2 Purely imaginary eigenvalue case
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
    $l_{\rho}^{(c)}$ $=$ $\widetilde{\Phi}_{\rho}(t)\widetilde{f}^{(c)}(t)=\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]q^{(c)}+\sum_{k=1}^{m}\Phi_{k}l_{k\rho}^{(c)}(t)+O(t)$
@@
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\mu q^{T}&\ldots&(t^{1/\rho}\mu)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
    $l_{k\rho}^{(c)}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\mu)^{\rho-1}(q^{(c)})^{\ast}&\ldots&t^{1/\rho}\omega(q^{(c)})^{\ast}&0\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k>\rho.\end{array}\right$.
@@ 4.3 A special case
$H=\widehat{U}\left[\begin{array}[]{cc|cc}\imath\alpha I_{p}+N&0&G&0\\ 0&F_{2}&0&G_{2}\\ \hline\cr 0&0&-(\imath\alpha I_{p}+N)^{\ast}&0\\ 0&R_{2}&0&-F_{2}^{\ast}\end{array}\right]\widehat{U}^{\ast},\quad T=J_{n}K,$
where $\alpha\in{\mathbb{R}}$, $\widehat{U}$ is unitary and symplectic ($\widehat{U}^{\ast}J_{n}\widehat{U}=J_{n}$), $H_{C}:=\left[\begin{array}[]{cc}F_{2}&G_{2}\\ R_{2}&-F_{2}^{\ast}\end{array}\right]$ is Hamiltonian and $\imath\alpha\notin\Lambda(H_{C})$. $K$ is positive semidefinite, NN is defined in (1), and
@@
After a simple two-sided block permutation, we have a unitary matrix $U=\left[\begin{array}[]{cc}U_{T}&U_{C}\end{array}\right]$ such that
$HU=U\left[\begin{array}[]{cc}H_{T}&0\\ 0&H_{C}\end{array}\right],\quad U^{\ast}JU=\left[\begin{array}[]{cc}\widehat{J}_{p}&0\\ 0&J_{n-p}\end{array}\right],$
@@
$H_{T}=\mathop{\rm diag}\nolimits(\imath\alpha I_{2s_{1}}+\widehat{H}_{2},\imath\alpha I_{4s_{2}}+\widehat{H}_{4},\ldots,\imath\alpha I_{2ms_{m}}+\widehat{H}_{2m}),\quad\widehat{J}_{p}=\mathop{\rm diag}\nolimits(J_{s_{1}},J_{2s_{2}},\ldots,J_{ms_{m}}),$
@@
$\Pi_{T}=\mathop{\rm diag}\nolimits(\Pi_{1},\ldots,\Pi_{m}).$
with
@@
One can always choose $Q$ such that $\Delta_{\Omega}=I_{\ell}$. When $t>0$ is sufficiently small, following Theorem 13 there is $\widetilde{\Phi}_{\Omega}(t)$ and
@@
Since the left hand side is Hermitian, so is $\widetilde{\Omega}(t)$. This implies that $H+tJ_{n}K$ has $\ell$ (semi)simple purely imaginary eigenvalues of the form $\imath(\alpha+t^{1/(2\rho)}\mu_{j,+}+o(t^{1/(2\rho)}))$ and the corresponding structure inertia indices are all $1$. In the same way, one can show that $H+tJ_{n}K$ has $\ell$ purely imaginary eigenvalues of the form $\imath(\alpha+t^{1/(2\rho)}\mu_{j,-}+o(t^{1/(2\rho)}))$ and the corresponding structure inertia indices are all 1-1. The same result was derived in [22], which now can be considered as a special case of Theorem 13.

### v2

@@ 4.2 Purely imaginary eigenvalue case
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
    $l_{\rho}^{(c)}(t)$ $=$ $\widetilde{\Phi}_{\rho}(t)\widetilde{f}^{(c)}(t)=\Upsilon_{\rho}\left[\begin{array}[]{cc}I_{s_{\rho}}\\ G_{\rho}\end{array}\right]q^{(c)}+\sum_{k=1}^{m}\Phi_{k}l_{k\rho}^{(c)}(t)+O(t)$
@@
    $l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\mu q^{T}&\ldots&(t^{1/\rho}\mu)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
    $l_{k\rho}^{(c)}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\mu)^{\rho-1}(q^{(c)})^{\ast}&\ldots&t^{1/\rho}\omega(q^{(c)})^{\ast}&0\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
@@ 4.3 A special case
$H=\widehat{U}\left[\begin{array}[]{cc|cc}\imath\alpha I_{p}+N&0&G&0\\ 0&F_{2}&0&G_{2}\\ \hline\cr 0&0&-(\imath\alpha I_{p}+N)^{\ast}&0\\ 0&R_{2}&0&-F_{2}^{\ast}\end{array}\right]\widehat{U}^{-1},\quad T=J_{n}K,$
where $\alpha\in{\mathbb{R}}$, $\widehat{U}$ is symplectic ($\widehat{U}^{\ast}J_{n}\widehat{U}=J_{n}$), $H_{C}:=\left[\begin{array}[]{cc}F_{2}&G_{2}\\ R_{2}&-F_{2}^{\ast}\end{array}\right]$ is Hamiltonian and $\imath\alpha\notin\Lambda(H_{C})$. $K$ is positive semidefinite, NN is defined in (1), and
@@
After a simple two-sided block permutation, we have a nonsingular matrix $U=\left[\begin{array}[]{cc}U_{T}&U_{C}\end{array}\right]$ such that
$HU=U\left[\begin{array}[]{cc}H_{T}&0\\ 0&H_{C}\end{array}\right],\quad U^{\ast}J_{n}U=\left[\begin{array}[]{cc}\widehat{J}_{p}&0\\ 0&J_{n-p}\end{array}\right],$
@@
$H_{T}=\mathop{\rm diag}\nolimits(\imath\alpha I_{2s_{1}}+\widehat{H}_{2},\imath\alpha I_{4s_{2}}+\widehat{H}_{4},\ldots,\imath\alpha I_{2ms_{m}}+\widehat{H}_{2m}),\quad\widehat{J}_{p}=\mathop{\rm diag}\nolimits(J_{s_{1}},J_{2s_{2}},\ldots,J_{ms_{m}})$
@@
$\Pi_{T}=\mathop{\rm diag}\nolimits(\Pi_{1},\ldots,\Pi_{m}),$
where
@@
One can always choose $Q$ such that $\Delta_{\Omega}=I_{\ell}$. When $t>0$ is sufficiently small, following Theorem 15 there are $\widetilde{\Phi}_{\Omega}(t)$ and
@@
Since the left-hand side is Hermitian, so is $\widetilde{\Omega}(t)$. This implies that $H+tJ_{n}K$ has $\ell$ (semi)simple purely imaginary eigenvalues of the form $\imath(\alpha+t^{1/(2\rho)}\mu_{j,+}+o(t^{1/(2\rho)}))$ and the corresponding structure inertia indices are all $1$. In the same way, one can show that $H+tJ_{n}K$ has $\ell$ purely imaginary eigenvalues of the form $\imath(\alpha+t^{1/(2\rho)}\mu_{j,-}+o(t^{1/(2\rho)}))$ and the corresponding structure inertia indices are all 1-1. The same result was derived in [22], which now can be considered as a special case of Theorem 15.
@@ 5 Conclusions
Acknowledgement The author thanks the anonymous referee and the handling editor for their comments and suggestions.